### v1

@@ Abstract
When two stiff inclusions are closely located, the gradient of the solution may become arbitrarily large as the distance between two inclusions tends to zero. Since blow-up of the gradient occurs in the narrow region, fine meshes should be required to compute the gradient. Thus, it is a challenging problem to numerically compute the gradient. Recent studies have shown that the major singularity can be extracted in an explicit way, so it suffices to compute the residual term for which only regular meshes are required. In this paper, we show through numerical simulations that the characterization of the singular term method can be efficiently used for the computation of the gradient when two strongly convex stiff domains of general shapes are closely located.
@@ 1 Introduction
Let $D_{1}$ and $D_{2}$ be two closely located strictly convex simply connected domains in $\mathbb{R}^{2}$ with $C^{2,\gamma}$ smooth boundaries for some $\gamma\in(0,1)$, see Figure 1. Let
@@
Let $H$ be a given entire harmonic function in $\mathbb{R}^{2}$. We consider the following problem
@@
Here and throughout this paper, $\nu$ denotes the unit outward normal on $\partial D_{j}$. The notations $|_{+}$ and $|_{-}$ are for limits from outside and inside of inclusions, respectively. It is worth mentioning that the constants $\lambda_{1}$ and $\lambda_{2}$ may or may not be the same depending on the applied field $H$.
@@
When $D_{1}$ and $D_{2}$ are closely located, the gradient of the solution $u$ may become arbitrarily large (blow-up) as the distance between two inclusions tends to zero. Two inclusions $D_{1}$ and $D_{2}$ may represent two perfect conductors of infinite conductivity embedded in a relatively weak conducting matrix. The solution $u$ represents the electric potential and the gradient of it represents the electric field. In two dimension, $D_{1}$ and $D_{2}$ may also represent the two dimensional cross-sections of two parallel elastic fibers embedded in an infinite elastic matrix. In this case, the solution $u$ represents the out-of-plane elastic displacement, and the gradient of the solution is proportional to the shear stress. When the inclusions are fiber-reinforced composites that are densely packed, the stress concentration may occur and cause material failure due to the damage of fiber composites. Thus, it is important to quantitatively understand the stress concentration. This problem was first raised in [6]. During the last two decades or more, significant development on this problem has been developed. It has been proved that the gradient blow-up rate is $\epsilon^{-1/2}$ in two dimensions [3, 4, 5, 7, 9, 11, 12, 18, 23, 24, 25], and $|\epsilon\ln\epsilon|^{-1}$ in three dimensions [8, 9, 10, 16, 19, 20, 21, 22], see [13] for more references.
@@
Due to the stress concentration near the narrow region between two inclusions, fine meshes are required to numerically compute the stress. Recently, a better understanding of the stress concentration has been proposed in [24] and used for various different circumstances [1, 14, 15, 16, 17]. It is shown that the asymptotic behavior of the gradient of the solution can be characterized by a singular function associated with two inclusions as the distance $\epsilon$ tends to zero. Using this singular function, the solution can be decomposed into a singular and a regular term. After extracting the singular term in an explicit way, it is sufficient to compute the residual term only using regular meshes. Thus, this characterization of the singular term method should have good applications in the numerical computation of the gradient of the solution. In fact, this idea was exploited numerically in [14] for the special case when two inclusions are disks. The numerical results show that the characterization of the singular term method can be effectively used for computation of the gradient in the presence of two nearly touching disks.
@@
Motivated by the theoretical result of characterization of the singular term method and its significant implication on numerical computation of the stress, in this paper, we numerically compute the solution for two nearly touching inclusions of general shapes and show the convergence of the solution. The main difference between the computation for two general shaped nearly touching inclusions and that in [14] for disks lies in the computation of the so called stress concentration factor which is the normalized magnitude of the stress concentration, as well as in the computation of the singular function. Those two terms are explicit and of simple form for two nearly touching disks, but they are not for general shapes. In this paper, we will show how to numerically obtain these two terms for nearly touching inclusions of general shapes. In fact, it was shown in [15] that the stress concentration factor converges to a certain integral of the solution to the touching case as the distance between two inclusions tends to zero. Based on this theoretical result, we will compute the value of the stress concentration factor accurately by numerically solving the touching case. The main goal of this paper is to show through numerical simulations that the characterization of the singular term method can be efficiently used for computation of the gradient when two closely located inclusions are of general shapes.
@@ 2 Review on the characterization of the singular term method
In this section, we briefly review the characterization of the singular term method obtained in [1, 15].
@@
Since $\nabla q$ blows up at the order of $\epsilon^{-1/2}$, $\alpha_{0}\nabla q$ is the singular part of $\nabla u$. Here, $\alpha_{0}$ is the so called stress concentration factor, which is given by the solution to the touching case, namely, the case when $\epsilon=0$. In fact, it is shown in [15] that $\alpha_{0}$ can be computed in the following way. For $\rho>0$, let
@@
By (2.6), one can obtain an accurate approximation of the stress concentration factor $\alpha_{0}$ by computing (2.5) through the touching case (2.4).
In particularly, if two inclusions are disks of radius $r_{1}$ and $r_{2}$, respectively, the stress concentration factor is given as
$\alpha_{0}=\frac{4\pi r_{1}r_{2}}{r_{1}+r_{2}}(\nu\cdot\nabla H)(z).$
However, the concentration factor could not be obtained explicitly for general shaped domain. Thanks to [3], the stress concentration factor was shown to converge to a certain integral of the solution to the touching case as the distance between two inclusions tends to zero. Based on this theoretical result, we will show how to compute the value of the stress concentration factor accurately by numerically solving the touching case in the next section.
Therefore, by (2.2), the solution $u$ to the problem (1.1) can be decomposed as a singular term and a regular term:
@@
for a constant $C$ independent of $\epsilon$. Therefore, for the numerical computation, it is sufficient to compute $b$ only using regular meshes.
@@ 6 Conclusion
In this paper, we show through numerical simulations that the computation of the stress concentration between closely located stiff inclusions of general shapes can be realized by using only regular meshes. Using the characterization of the singular term method, we can decompose the solution into a singular and a regular term. After extracting the singular in a precise way, we can compute the remaining term using regular meshes. The key point in our computation lies in the computation of the stress concentration factor as well as the singular term. We have shown that the computation of the stress concentration factor converges very fast. By comparing the convergent rate with the solution computed using layer potential techniques in a direct way, we conclude that the characterization of the singular term method can be used effectively for the computation of the solution.
Acknowledgement
The authors would like to express their gratitude to Hyeonbae Kang for pointing out this work as well as his kind advice.

### v2

@@ Abstract
When two stiff inclusions are closely located, the gradient of the solution to the Lamé system, in other words the stress, may become arbitrarily large as the distance between two inclusions tends to zero. To compute the gradient of the solution in the narrow region, extremely fine meshes are required. It is a challenging problem to numerically compute the stress near the narrow region between two inclusions of general shapes as their distance goes to zero. A recent study [15] has shown that the major singularity of the gradient can be extracted in an explicit way for two general shaped inclusions. Thus the complexity of the computation can be greatly reduced by removing the singular term and it suffices to compute the residual term only using regular meshes. The goal of this paper is to numerically compute the stress concentration in a fast and efficient way. In this paper, we compute the value of the stress concentration factor, which is the normalized magnitude of the stress concentration, for general shaped domain as the distance between two inclusions tends to zero. We also compute the solution for two closely located inclusions of general shapes and show the convergence of the solution. Only regular meshes are used in our numerical computation and the results clearly show that the characterization of the singular term method can be efficiently used for computation of the stress concentration between two closely located inclusions of general shapes.
@@ 1 Introduction
Let $D_{1}$ and $D_{2}$ be two closely located strictly convex simply connected domains in $\mathbb{R}^{2}$ with $C^{2,\gamma}$ smooth boundaries for some $\gamma\in(0,1)$, see, for example, Figure 1. Let
@@
Let $H$ be a given entire harmonic function in $\mathbb{R}^{2}$. We consider the following two dimensional problem
@@
Here and throughout this paper, $\nu$ denotes the unit outward normal on $\partial D_{j}$. The notations $|_{+}$ and $|_{-}$ are for limits from outside and inside of inclusions, respectively. It is worth mentioning that the constants $\lambda_{1}$ and $\lambda_{2}$ may or may not be the same depending on the applied field $H$. Two inclusions $D_{1}$ and $D_{2}$ may represent the two dimensional cross-sections of two parallel elastic fibers embedded in an infinite elastic matrix. The solution $u$ represents the out-of-plane elastic displacement, and the gradient of the solution is proportional to the shear stress. When the inclusions are fiber-reinforced composites that are densely packed, the stress concentration may occur and cause material failure due to the damage of fiber composites. Thus, it is important to quantitatively understand the stress concentration. This problem was first raised in [6]. During the last two decades or more, significant development on this problem has been developed. It has been proved that the gradient blow-up rate is $\epsilon^{-1/2}$ in two dimensions [3, 4, 5, 7, 9, 11, 12, 18, 23, 24, 25], and $|\epsilon\ln\epsilon|^{-1}$ in three dimensions [8, 9, 10, 16, 19, 20, 21, 22], see [13] for more references.
@@
As for the numerical computation, due to the stress concentration near the narrow region between two inclusions, extremely fine meshes are required to numerically compute the stress. Recently, a better understanding of the stress concentration has been proposed in [24] and used for various different circumstances [1, 14, 15, 16, 17]. It is shown that the asymptotic behavior of the gradient of the solution can be characterized by a singular function associated with two inclusions as the distance $\epsilon$ tends to zero. Using this singular function, the solution can be decomposed into a singular and a regular term. After extracting the singular term in an explicit way, it is sufficient to compute the residual term only using regular meshes. Therefore, the complexity of computation can be greatly reduced by removing the singular term. In fact, this idea was exploited numerically in [14] for the special case when two inclusions are of circular shapes. The stress concentration factor, which is the normalized magnitude of the stress concentration, is only related to the radii of two disks as well as the applied field which is easy to compute. Meanwhile, the singular function which characterizes the singular behavior of the stress is also explicit for circular domains. The numerical results in [14] showed that the characterization of the singular term method can be effectively used for computation of the gradient in the presence of two nearly touching disks.
@@
However, if two nearly touching inclusions are of general shapes, it is difficult to find accurate stress concentration factor since it is related to the global information of the singular function which is not explicit for general shapes. Thanks to the study in [15], where it shows that as the distance between two inclusions tends to zero, the stress concentration factor converges to a certain integral of the solution to the system when two inclusions are touching. Our goal of this paper is to numerically compute the stress concentration between two closely located stiff inclusions of general shapes in a fast and accurate way. Motivated by the idea of characterization of the singular term and the fact that the stress concentration factor converges to a certain integral of the solution to the touching case which is shown in [15], we are able to obtain fast and accurate numerical computation of the stress concentration only using regular meshes. We will compute the value of the stress concentration factor accurately by numerically solving the touching case. We then numerically compute the solution for two closely located inclusions of general shapes and show the convergence of the solution. Our numerical computations clearly show that the characterization of the singular term method can be efficiently used for computation of the stress concentration when two closely located inclusions are of general shapes. It is worth mentioning that only regular meshes are used in our computations.
@@ 2 Review on the characterization of the singular term method
In this section, we briefly review the characterization of the singular term method in [1, 15].
@@
Since $\nabla q$ blows up at the order of $\epsilon^{-1/2}$, $\alpha_{0}\nabla q$ is the singular part of $\nabla u$. Here, $\alpha_{0}$ is the so called stress concentration factor. In particularly, if two inclusions are disks of radii $r_{1}$ and $r_{2}$, respectively, the stress concentration factor is given as
$\alpha_{0}=\frac{4\pi r_{1}r_{2}}{r_{1}+r_{2}}(\nu\cdot\nabla H)(z).$
For general shaped inclusions, the stress concentration factor was shown, in [15], to converge to a certain integral of the solution to the touching case as the distance between two inclusions tends to zero, namely, the case when $\epsilon=0$. In fact, it is shown in [15] that $\alpha_{0}$ can be computed in the following way. For $\rho>0$, let
@@
By (2.6), one can obtain an accurate approximation of the stress concentration factor $\alpha_{0}$ by computing (2.5) through the touching case (2.4). Therefore, by (2.2), the solution $u$ to the problem (1.1) can be decomposed as a singular term and a regular term:
@@
for a constant $C$ independent of $\epsilon$. For the numerical computation, it is sufficient to compute $b$ only using regular meshes. As it is shown above, it is a key point to obtain accurate value of the stress concentration factor $\alpha_{0}$. In the next section, we will show how to obtain the accurate value of the stress concentration factor through numerically solving the touching case.
@@ 6 Conclusion
In this paper, we show through numerical simulations that the computation of the stress concentration between closely located stiff inclusions of general shapes can be realized by only using regular meshes. Using the characterization of the singular term method, we can decompose the solution into a singular and a regular term. After extracting the singular in a precise way, we can compute the remaining term using regular meshes. The key point in our computation lies in the computation of the stress concentration factor as well as the singular term. We have shown that the computation of the stress concentration factor converges very fast. By comparing the convergent rate with the solution computed using layer potential techniques in a direct way, we conclude that the characterization of the singular term method can be used effectively for the computation of the solution. This numerical method could also be generalized to the three dimensional case, which we will investigate in a forthcoming work.